OpenStudy (anonymous):

use the slope and y intercept method graph the following equatios 3x-4y-1=27

OpenStudy (saifoo.khan):

lol, same? again?1

OpenStudy (saifoo.khan):

simple thing.. always remember while graphing. when x intercpts y = 0 when y intercpts x = 0. Insert to get the intercepts respectively.

myininaya (myininaya):

3x-4y-1=27 3x-4y-1+1=27+1 3x-4y=28 3x-3x-4y=28-3x -4y=-3x+28 y=-3/(-4) *x+28/(-4) y=3/4 *x-7 y intercept is -7 (so a draw a point on the y axis at -7) then you 3/4 to find another point from that point using -7 on the y axis count up 3 units from that then go to the right 4 units and put another dot then connect the dots

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (saifoo.khan):

She showed the whole method/